# Integrating Dynamical Systems Modeling with Spatiotemporal scRNA-seq Data Analysis

### v1

@@ 3.2.2 Continuous Temporal Dynamics Modeling
It can be proved that when $\sigma\rightarrow 0$, this also gives a way to solve the dynamical optimal transport (Tong et al., 2024a). The advantage of CFM is that it is simulation-free and can handle the thousand gene dimensions without reducing the dimensionality.
Schrödinger Bridge Problem ($g=0,\sigma\neq 0$)
@@
In this case, the model can account for the stochastic effects yet without unbalanced effects. We employ the Schrödinger Bridge problem to model the SDE dynamics, i,e, the cellular dynamics is $\mathrm{d}{\mathbf{x}}_{t}=\mathbf{b}\left({\mathbf{x}}_{t},t\right)\mathrm{d}t+\boldsymbol{\sigma}\left({\mathbf{x}}_{t},t\right)\mathrm{d}\mathbf{w}_{t}$. The Schrödinger Bridge problem seeks to determine the most probable evolution between a specified initial distribution $\nu_{0}$ and a terminal distribution $\nu_{1}$ (assumed to possess a density in this study) relative to a given reference stochastic process. Formally, this problem is formulated as the minimization of the Kullback-Leibler (KL) divergence from the perspective of optimal control (Dai Pra, 1991), as shown below:
$\min_{\mu^{\mathbf{X}}_{0}=\nu_{0},\mu^{\mathbf{X}}_{1}=\nu_{1}}\mathcal{D}_{\mathrm{KL}}\left(\mu^{\mathbf{X}}_{[0,1]}|\mu^{\mathbf{Y}}_{[0,1]}\right),$ (3.11)
@@
where $\mu^{\mathbf{X}}_{[0,1]}$ denotes the probability measure induced by the stochastic process ${\mathbf{x}}_{t}$ for $0\leq t\leq 1$, defined on the space of all continuous paths $C([0,1],\mathbb{R}^{d})$. The distribution of ${\mathbf{x}}_{t}$ at a given time t is characterized by the measure $\mu^{\mathbf{X}}_{t}$ with density function $p(\mathbf{x},t)$. The reference measure $\mu^{\mathbf{Y}}_{[0,1]}$ is chosen as the probability measure induced by the process $\mathrm{d}\mathbf{Y}_{t}=\boldsymbol{\sigma}(\mathbf{Y}_{t},t)\mathrm{d}\mathbf{w}_{t},$ where $\mathbf{w}_{t}\in\mathbb{R}^{d}$ represents the standard multidimensional Brownian motion.
@@
Interestingly, the problem can be equivalently transformed into a dynamical form (Gentil et al., 2017; Chen et al., 2016; Dai Pra, 1991; Zhang et al., 2025)
$\inf_{\left(p,\mathbf{b}\right)}\int_{0}^{1}\int_{\mathbb{R}^{d}}\left[\frac{1}{2}\mathbf{b}^{\top}(\mathbf{x},t)\mathbf{a}(\mathbf{x},t)^{-1}\mathbf{b}(\mathbf{x},t)\right]p(\mathbf{x},t)\mathrm{d}\mathbf{x}\mathrm{d}t,$ (3.12)
@@
where the infimum is taken over all pairs of functions $(p,\mathbf{b})$ satisfying $p(\cdot,0)=\nu_{0}$, $p(\cdot,1)=\nu_{1}$, and $p(\mathbf{x},t)$ is absolutely continuous with respect to time. Additionally, the pair $(p,\mathbf{b})$ must satisfy the Fokker-Planck Equation (2.3). We denote minimization problem (3.12) and the constraints (2.3) as the dynamic diffusion Schrödinger bridge formulation. Methods for modeling stochastic dynamics based on it have been widely developed (Maddu et al., 2024; Zhou et al., 2024b; Jiang and Wan, 2024; Yeo et al., 2021; Koshizuka and Sato, 2023; Tong et al., 2024b), involving neural SDE, neural ODE, or flow matching techniques. We will next provide an overview of the methodologies in these approaches.
Nerual SDE Solver
@@
$\inf_{({p}(\mathbf{x},t),\mathbf{b}(\mathbf{x},t))}\mathbb{E}_{\mathbf{x}_{0}\sim\nu_{0}}\int_{0}^{1}\left[\frac{1}{2}\mathbf{b}^{\top}(\mathbf{x}(t),t)\mathbf{a}(\mathbf{x}(t),t)^{-1}\mathbf{b}(\mathbf{x}(t),t)\right]\mathrm{d}t+\lambda\mathcal{D}(p(\mathbf{x},1),\nu_{1}).$
Building upon this one can parametrize $\mathbf{b}(\mathbf{x},t)$ and $\boldsymbol{\sigma}(\mathbf{x},t)$ using neural networks respectively and solve this formulation through POT and the neural SDE solver. However, besides these two terms, some work also introduces the idea of the principle of least action along the trajectory in which the optimal path has the smallest action value (Jiang and Wan, 2024; Koshizuka and Sato, 2023). Thus they introduce a new Hamilton–Jacobi–Bellman (HJB) regularization term (Jiang and Wan, 2024) when assuming $\mathbf{b}(\mathbf{x},t)=-\nabla_{\mathbf{x}}\Phi(\mathbf{x},t)$, i.e.,
$\mathcal{R}_{h}=\int_{0}^{1}\int_{\mathbb{R}^{d}}\left|\partial_{t}\Phi(\mathbf{x},t)-\|\nabla_{\mathbf{x}}\Phi(\mathbf{x},t)\|_{2}^{2}\right|{p}(\mathbf{x},t)\mathrm{d}\mathbf{x}\mathrm{d}t.$
or a general form derived in (Koshizuka and Sato, 2023).
@@
By leveraging CFM techniques, the simulation-free Shrödinger bridge (Tong et al., 2024b) has also been recently developed. The core idea is to decompose the problem into a sequence of elementary conditional subproblems, each of which is more tractable, and subsequently express the overall solution as a mixture of the solutions to these conditional subproblems. Let the reference process be a Brownian motion (i.e., $\mathbf{Y}=\sigma\mathbf{W}$). In this case, the Schrödinger bridge problem admits a unique solution $\mathbb{P}^{*}$, which is expressed as a mixture of Brownian bridges weighted by an entropic optimal transport (OT) plan:
$\mathbb{P}^{*}\left(\left(\mathbf{x}_{t}\right)_{t\in[0,1]}\right)=\int\mathbb{W}\left(\left(\mathbf{x}_{t}\right)\mid\mathbf{x}_{0},\mathbf{x}_{1}\right)\,\mathrm{d}\pi_{2\sigma^{2}}^{\star}\left(x_{0},x_{1}\right),$ (3.13)
where $\mathbb{W}\left(\left(\mathbf{x}_{t}\right)\mid{t\in(0,1)}\mid\mathbf{x}_{0},\mathbf{x}_{1}\right)$ denotes the Brownian bridge between $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$ with a diffusion rate $\sigma$, and $\pi_{2\sigma^{2}}^{\star}(\mathbf{x}_{0},\mathbf{x}_{1})$ represents the entropic optimal transport plan between the distributions. The calculation of $\mathbb{W}\left(\left(\mathbf{x}_{t}\right)\mid{t\in(0,1)}\mid\mathbf{x}_{0},\mathbf{x}_{1}\right)$ can be framed as an optimal control problem:
@@
$\mathrm{d}{\mathbf{x}}_{t}=\mathbf{b}\left({\mathbf{x}}_{t},t\right)\mathrm{d}t+\boldsymbol{\sigma}\mathrm{d}\mathbf{w}_{t},$
$\mathbf{X}_{0}\sim\delta_{x_{0}},\quad\mathbf{X}_{1}\sim\delta_{x_{1}},$
@@
Assume $\sigma$ is constant and then the corresponding Fokker-Planck equation in (3.11) yields:
$\partial_{t}p(\mathbf{x},t)=-\nabla_{\mathbf{x}}\cdot\left(p(\mathbf{x},t)\mathbf{b}(\mathbf{x},t)\right)+\frac{1}{2}\sigma^{2}\Delta p(\mathbf{x},t).$
@@
$\mathrm{d}\mathbf{X}_{t}=\underbrace{\left(\mathbf{b}\left(\mathbf{X}_{t},t\right)-\frac{1}{2}\sigma^{2}\nabla_{\mathbf{x}}\log p\left(\mathbf{X}_{t},t\right)\right)}_{v(\mathbf{X}_{t},t)}\mathrm{d}t,$ (3.14)
together with the initial distribution generate the same distribution as SDE. The (3.14) is called the probability flow ODE. Conversely, if the probability flow ODE $\mathbf{v}(\mathbf{x},t)$ and $\nabla_{\mathbf{x}}\log p(\mathbf{x},t)$ (also known as score function) are known, one can recover the SDE drift through $\mathbf{v}(\mathbf{x},t)=\mathbf{b}(\mathbf{x},t)+\frac{1}{2}\sigma^{2}\nabla_{\mathbf{x}}\log{p}(\mathbf{x},t)$. So the flow-matching objective is
$\mathcal{L}_{\mathrm{U}[\mathrm{SF}]^{2}\mathrm{M}}(\theta)=\mathbb{E}[\underbrace{\left\|\mathbf{v}_{\theta}(\mathbf{x},t)-\mathbf{v}(\mathbf{x},t)\right\|^{2}}_{\text{flow matching loss }}+\lambda(t)^{2}\underbrace{\left\|\nabla s_{\theta}(\mathbf{x},t)-\nabla\log p(\mathbf{x},t)\right\|^{2}}_{\text{score matching loss }}].$
However, this loss is intractable, by (3.13) and the CFM objective, one can transform it into a trackable loss
$\mathbb{E}_{Q^{\prime}}\underbrace{\left\|\mathbf{v}_{\theta}(\mathbf{x},t)-\mathbf{v}(\mathbf{x},t\mid(\mathbf{x}_{0},\mathbf{x}_{1})\right\|^{2}}_{\text{conditional flow matching loss }}\\ +\mathbb{E}_{Q^{\prime}}\lambda(t)^{2}\underbrace{\left\|\nabla s_{\theta}(\mathbf{x},t)-\nabla\log p(\mathbf{x},t\mid(\mathbf{x}_{0},\mathbf{x}_{1})\right\|^{2}}_{\text{conditional score matching loss }},$
where $Q^{\prime}=t\sim\mathcal{U}(0,1)\otimes q(\mathbf{x}_{0},\mathbf{x}_{1})\otimes p(\mathbf{x},t|(\mathbf{x}_{0},\mathbf{x}_{1}))$. Since the conditional path is a Brownian bridge, the analytic form can be derived, i.e., $p(\mathbf{x},t\mid(\mathbf{x}_{0},\mathbf{x}_{1}))=\mathcal{N}\left(\mathbf{x};t\mathbf{x}_{1}+(1-t)\mathbf{x}_{0},\sigma^{2}t(1-t)\right)$ and
$\mathbf{v}(\mathbf{x},t\mid(\mathbf{x}_{0},\mathbf{x}_{1}))=\frac{1-2t}{t(1-t)}\left(\mathbf{x}-\left(t\mathbf{x}_{1}+(1-t)\mathbf{x}_{0}\right)\right)+\left(\mathbf{x}_{1}-\mathbf{x}_{0}\right),$
$\nabla_{\mathbf{x}}\log p(\mathbf{x},t\mid(\mathbf{x}_{0},\mathbf{x}_{1}))=\frac{t\mathbf{x}_{1}+(1-t)\mathbf{x}_{0}-\mathbf{x}}{\sigma^{2}t(1-t)},~{}t\in[0,1].$
And $q(\mathbf{x}_{0},\mathbf{x}_{1})$ can be computed by the entropic optimal transport.
Unbalanced Wasserstein-Fisher-Rao metric ($g\neq 0,\sigma=0$)
In this case, the model can account for the unbalanced dynamics, however, it can not account for stochastic dynamics. The cellular dynamics is also governed by the ODE model by $\mathrm{d}{\mathbf{x}}_{t}=\mathbf{b}({\mathbf{x}}_{t},t)\mathrm{d}t$. Then one can use the dynamical unbalanced optimal transport to model these dynamics also known as Wasserstein-Fisher-Rao metric (Chizat et al., 2018a, b; Gangbo et al., 2019), which can be stated as
@@
Here $\alpha$ denotes a hyperparameter that controls the weighting. It is also important to note that in this context, $\nu_{0}$ and $\nu_{1}$ do not necessarily correspond to normalized probability densities; rather, they generally represent mass densities.
@@
Recent works such as TrajectoryNet and TIGON utilize (3.15) to infer unbalanced dynamics from scRNA-seq data (Sha et al., 2024; Tong et al., 2023). To derive a CNF solver for (3.15), TIGON(Sha et al., 2024) observes that along the characteristic line $\frac{\mathrm{d}\mathbf{x}}{\mathrm{d}t}=\mathbf{b}(\mathbf{x},t)$, one have
$\int_{0}^{1}\int_{\mathbb{R}^{d}}f(\mathbf{x},t)p(\mathbf{x},t)\mathrm{d}\mathbf{x}\mathrm{d}t=\mathbb{E}_{\mathbf{x_{0}}\sim p_{0}}\int_{0}^{1}f(\mathbf{x},t)e^{\int_{0}^{t}g(\mathbf{x},s)\mathrm{d}s}\mathrm{d}t$
@@
and $\frac{\mathrm{d}(\ln p)}{\mathrm{d}t}=g-\nabla\cdot\mathbf{v}$. This can make the computation of both energy loss and reconstruction loss in high dimensional space tractable. Therefore, one can parameterize $\mathbf{b}(\mathbf{x},t)$ and $g(\mathbf{x},t)$ using neural networks respectively, and train them by minimizing the overall loss.
Regularized Unbalanced Optimal Transport ($g\neq 0,\sigma\neq 0$)
@@
In this case, the model can account for both the unbalanced and stochastic dynamics. The cellular dynamics is governed by the SDE model $\mathrm{d}{\mathbf{x}}_{t}=\mathbf{b}\left({\mathbf{x}}_{t},t\right)\mathrm{d}t+\sigma(t)\mathbf{I}\mathrm{d}\mathbf{w}_{t}$. We can use the regularized unbalanced optimal transport to model it (Zhang et al., 2025; Baradat and Lavenant, 2021). It can be viewed as an unbalanced relaxation of the dynamic formulation of Schrödinger bridge problem. Consider
$\inf_{\left(p,\mathbf{b},g\right)}\int_{0}^{1}\int_{\mathbb{R}^{d}}\frac{1}{2}\left\|\mathbf{b}(\mathbf{x},t)\right\|_{2}^{2}{p}(\mathbf{x},t)\mathrm{d}\mathbf{x}\mathrm{d}t+\int_{0}^{1}\int_{\mathbb{R}^{d}}\alpha\Psi\left(g(\mathbf{x},t)\right){p}(\mathbf{x},t)\mathrm{d}\mathbf{x}\mathrm{d}t,$ (3.16)
@@
where $\Psi:\mathbb{R}\rightarrow[0,+\infty]$ corresponds to the growth penalty function, the infimum is taken over all pairs $\left(p,\mathbf{b}\right)$ such that $p(\cdot,0)=$ $\nu_{0},p(\cdot,1)=\nu_{1},p(\mathbf{x},t)$ absolutely continuous, and
@@
$\partial_{t}p=-\nabla_{\mathbf{x}}\cdot\left(p\mathbf{b}\right)+\frac{1}{2}\nabla_{\mathbf{x}}^{2}:\left(\sigma^{2}(t)\mathbf{I}p\right)+gp$ (3.17)
with vanishing boundary condition: $\lim_{|\mathbf{x}|\rightarrow\infty}p(\mathbf{x},t)=0$.
One can similarly develop a dynamical OT solver relying on a neural SDE solver, which might be less efficient compared to a neural ODE solver. Recently, DeepRUOT (Zhang et al., 2025)reformulates the RUOT problem with the Fisher information regularization, equivalently expressed as
@@
$\inf_{\left(p,\mathbf{v},g\right)}\int_{0}^{1}\int_{\mathbb{R}^{d}}\left[\frac{1}{2}\left\|\mathbf{v}(\mathbf{x},t)\right\|_{2}^{2}+\frac{\sigma^{4}(t)}{8}\left\|\nabla_{\mathbf{x}}\log p\right\|_{2}^{2}-\frac{\sigma^{2}(t)}{2}\left(1+\log p\right)g+\alpha\Psi\left(g\right)\right]{p}(\mathbf{x},t)\mathrm{d}\mathbf{x}\mathrm{d}t,$ (3.18)

### v2

@@ 3.2.2 Continuous Temporal Dynamics Modeling
It can be proved that when $\sigma\rightarrow 0$, this also gives a way to solve the dynamical optimal transport Tong et al. (2024). The advantage of CFM is that it is simulation-free and can handle the thousand gene dimensions without reducing dimensionality.
@@
Schrödinger Bridge Problem ($g=0,\sigma\neq 0$) In this case, the model can account for the stochastic effects, yet without unbalanced effects. We employ the Schrödinger Bridge problem to model the SDE dynamics, i.e., the cellular dynamics are $\mathrm{d}{\boldsymbol{x}}_{t}=\boldsymbol{b}\left({\boldsymbol{x}}_{t},t\right)\mathrm{d}t+\boldsymbol{\sigma}\left({\boldsymbol{x}}_{t},t\right)\mathrm{d}\boldsymbol{w}_{t}$. The Schrödinger Bridge problem seeks to determine the most probable evolution between a specified initial distribution $\nu_{0}$ and a terminal distribution $\nu_{1}$ (assumed to possess a density in this study) relative to a given reference stochastic process. Formally, this problem is formulated as the minimization of the Kullback–Leibler (KL) divergence from the perspective of optimal control Dai Pra (1991), as shown below:
$\min_{\mu^{\boldsymbol{X}}_{0}=\nu_{0},\mu^{\boldsymbol{X}}_{1}=\nu_{1}}\mathcal{D}_{\mathrm{KL}}\left(\mu^{\boldsymbol{X}}_{[0,1]}|\mu^{\boldsymbol{Y}}_{[0,1]}\right),$ (12)
@@
where $\mu^{\boldsymbol{X}}_{[0,1]}$ denotes the probability measure induced by the stochastic process ${\boldsymbol{x}}_{t}$ for $0\leq t\leq 1$, defined on the space of all continuous paths $C([0,1],\mathbb{R}^{d})$. The distribution of ${\boldsymbol{x}}_{t}$ at a given time t is characterized by the measure $\mu^{\boldsymbol{X}}_{t}$ with density function $p(\boldsymbol{x},t)$. The reference measure $\mu^{\boldsymbol{Y}}_{[0,1]}$ is chosen as the probability measure induced by the process $\mathrm{d}\boldsymbol{Y}_{t}=\boldsymbol{\sigma}(\boldsymbol{Y}_{t},t)\mathrm{d}\boldsymbol{w}_{t},$ where $\boldsymbol{w}_{t}\in\mathbb{R}^{d}$ represents the standard multidimensional Brownian motion.
@@
Interestingly, the problem can be equivalently transformed into a dynamical form Gentil et al. (2017); Chen et al. (2016); Dai Pra (1991); Zhang et al. (2025)
$\inf_{\left(p,\boldsymbol{b}\right)}\int_{0}^{1}\int_{\mathbb{R}^{d}}\left[\frac{1}{2}\boldsymbol{b}^{\top}(\boldsymbol{x},t)\boldsymbol{a}(\boldsymbol{x},t)^{-1}\boldsymbol{b}(\boldsymbol{x},t)\right]p(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t,$ (13)
@@
where the infimum is taken over all pairs of functions $(p,\boldsymbol{b})$ satisfying $p(\cdot,0)=\nu_{0}$, $p(\cdot,1)=\nu_{1}$, and $p(\boldsymbol{x},t)$ is absolutely continuous with respect to time. Additionally, the pair $(p,\boldsymbol{b})$ must satisfy the Fokker–Planck Equation (3). We denote minimization problem (13) and the constraints (3) as the dynamic diffusion Schrödinger bridge formulation. Methods for modeling stochastic dynamics based on it have been widely developed Maddu et al. (2024); Zhou et al. (2024); Jiang and Wan (2024); Yeo et al. (2021); Koshizuka and Sato (2023); Tong et al. (2024), involving neural SDE, neural ODE, or flow matching techniques. We will next provide an overview of the methodologies in these approaches.
Neural SDE Solver
@@
$\inf_{({p}(\boldsymbol{x},t),\boldsymbol{b}(\boldsymbol{x},t))}\mathbb{E}_{\boldsymbol{x}_{0}\sim\nu_{0}}\int_{0}^{1}\left[\frac{1}{2}\boldsymbol{b}^{\top}(\boldsymbol{x}(t),t)\boldsymbol{a}(\boldsymbol{x}(t),t)^{-1}\boldsymbol{b}(\boldsymbol{x}(t),t)\right]\mathrm{d}t+\lambda\mathcal{D}(p(\boldsymbol{x},1),\nu_{1}).$
Building upon this, one can parametrize $\boldsymbol{b}(\boldsymbol{x},t)$ and $\boldsymbol{\sigma}(\boldsymbol{x},t)$ using neural networks respectively and solve this formulation through POT and the neural SDE solver. However, besides these two terms, some work also introduces the idea of the principle of least action along the trajectory in which the optimal path has the smallest action value Jiang and Wan (2024); Koshizuka and Sato (2023). Thus, they introduce a new Hamilton–Jacobi–Bellman (HJB) regularization term Jiang and Wan (2024) when assuming $\boldsymbol{b}(\boldsymbol{x},t)=-\nabla_{\boldsymbol{x}}\Phi(\boldsymbol{x},t)$, i.e.,
$\mathcal{R}_{h}=\int_{0}^{1}\int_{\mathbb{R}^{d}}\left|\partial_{t}\Phi(\boldsymbol{x},t)-\|\nabla_{\boldsymbol{x}}\Phi(\boldsymbol{x},t)\|_{2}^{2}\right|{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t.$
or a general form derived in Koshizuka and Sato (2023).
@@
By leveraging CFM techniques, the simulation-free Shrödinger bridge Tong et al. (2024) has also been recently developed. The core idea is to decompose the problem into a sequence of elementary conditional subproblems, each of which is more tractable, and subsequently express the overall solution as a mixture of the solutions to these conditional subproblems. Let the reference process be a Brownian motion (i.e., $\boldsymbol{Y}=\sigma\boldsymbol{W}$). In this case, the Schrödinger bridge problem admits a unique solution $\mathbb{P}^{*}$, which is expressed as a mixture of Brownian bridges weighted by an entropic optimal transport (OT) plan:
$\mathbb{P}^{*}\left(\left(\boldsymbol{x}_{t}\right)_{t\in[0,1]}\right)=\int\mathbb{W}\left(\left(\boldsymbol{x}_{t}\right)\mid\boldsymbol{x}_{0},\boldsymbol{x}_{1}\right)\,\mathrm{d}\pi_{2\sigma^{2}}^{\star}\left(x_{0},x_{1}\right),$ (14)
where $\mathbb{W}\left(\left(\boldsymbol{x}_{t}\right)\mid{t\in(0,1)}\mid\boldsymbol{x}_{0},\boldsymbol{x}_{1}\right)$ denotes the Brownian bridge between $\boldsymbol{x}_{0}$ and $\boldsymbol{x}_{1}$ with a diffusion rate $\sigma$, and $\pi_{2\sigma^{2}}^{\star}(\boldsymbol{x}_{0},\boldsymbol{x}_{1})$ represents the entropic optimal transport plan between the distributions. The calculation of $\mathbb{W}\left(\left(\boldsymbol{x}_{t}\right)\mid{t\in(0,1)}\mid\boldsymbol{x}_{0},\boldsymbol{x}_{1}\right)$ can be framed as an optimal control problem:
@@
$\mathrm{d}{\boldsymbol{x}}_{t}=\boldsymbol{b}\left({\boldsymbol{x}}_{t},t\right)\mathrm{d}t+\boldsymbol{\sigma}\mathrm{d}\boldsymbol{w}_{t},$
$\boldsymbol{X}_{0}\sim\delta_{x_{0}},\quad\boldsymbol{X}_{1}\sim\delta_{x_{1}},$
@@
Assume $\sigma$ is constant and then the corresponding Fokker–Planck equation in (12) yields
$\partial_{t}p(\boldsymbol{x},t)=-\nabla_{\boldsymbol{x}}\cdot\left(p(\boldsymbol{x},t)\boldsymbol{b}(\boldsymbol{x},t)\right)+\frac{1}{2}\sigma^{2}\Delta p(\boldsymbol{x},t).$
@@
$\mathrm{d}\boldsymbol{X}_{t}=\underbrace{\left(\boldsymbol{b}\left(\boldsymbol{X}_{t},t\right)-\frac{1}{2}\sigma^{2}\nabla_{\boldsymbol{x}}\log p\left(\boldsymbol{X}_{t},t\right)\right)}_{v(\boldsymbol{X}_{t},t)}\mathrm{d}t,$ (15)
together with the initial distribution generate the same distribution as SDE. The (15) is called the probability flow ODE. Conversely, if the probability flow ODE $\boldsymbol{v}(\boldsymbol{x},t)$ and $\nabla_{\boldsymbol{x}}\log p(\boldsymbol{x},t)$ (also known as score function) are known, one can recover the SDE drift through $\boldsymbol{v}(\boldsymbol{x},t)=\boldsymbol{b}(\boldsymbol{x},t)+\frac{1}{2}\sigma^{2}\nabla_{\boldsymbol{x}}\log{p}(\boldsymbol{x},t)$. So the flow-matching objective is
$\mathcal{L}_{\mathrm{U}[\mathrm{SF}]^{2}\mathrm{M}}(\theta)=\mathbb{E}[\underbrace{\left\|\boldsymbol{v}_{\theta}(\boldsymbol{x},t)-\boldsymbol{v}(\boldsymbol{x},t)\right\|^{2}}_{\text{flow matching loss }}+\lambda(t)^{2}\underbrace{\left\|\nabla s_{\theta}(\boldsymbol{x},t)-\nabla\log p(\boldsymbol{x},t)\right\|^{2}}_{\text{score matching loss }}].$
However, this loss is intractable; by (14) and the CFM objective, one can transform it into a tractable loss
$\mathcal{L}_{[\mathrm{SF}]^{2}\mathrm{M}}(\theta)=\mathbb{E}_{Q^{\prime}}\underbrace{\left\|\boldsymbol{v}_{\theta}(\boldsymbol{x},t)-\boldsymbol{v}(\boldsymbol{x},t\mid(\boldsymbol{x}_{0},\boldsymbol{x}_{1})\right\|^{2}}_{\text{conditional flow matching loss }}\\ +\mathbb{E}_{Q^{\prime}}\lambda(t)^{2}\underbrace{\left\|\nabla s_{\theta}(\boldsymbol{x},t)-\nabla\log p(\boldsymbol{x},t\mid(\boldsymbol{x}_{0},\boldsymbol{x}_{1})\right\|^{2}}_{\text{conditional score matching loss }},$
where $Q^{\prime}=t\sim\mathcal{U}(0,1)\otimes q(\boldsymbol{x}_{0},\boldsymbol{x}_{1})\otimes p(\boldsymbol{x},t|(\boldsymbol{x}_{0},\boldsymbol{x}_{1}))$. Since the conditional path is a Brownian bridge, the analytic form can be derived, i.e., $p(\boldsymbol{x},t\mid(\boldsymbol{x}_{0},\boldsymbol{x}_{1}))=\mathcal{N}\left(\boldsymbol{x};t\boldsymbol{x}_{1}+(1-t)\boldsymbol{x}_{0},\sigma^{2}t(1-t)\right)$ and
$\boldsymbol{v}(\boldsymbol{x},t\mid(\boldsymbol{x}_{0},\boldsymbol{x}_{1}))=\frac{1-2t}{t(1-t)}\left(\boldsymbol{x}-\left(t\boldsymbol{x}_{1}+(1-t)\boldsymbol{x}_{0}\right)\right)+\left(\boldsymbol{x}_{1}-\boldsymbol{x}_{0}\right),$
$\nabla_{\boldsymbol{x}}\log p(\boldsymbol{x},t\mid(\boldsymbol{x}_{0},\boldsymbol{x}_{1}))=\frac{t\boldsymbol{x}_{1}+(1-t)\boldsymbol{x}_{0}-\boldsymbol{x}}{\sigma^{2}t(1-t)},~{}t\in[0,1].$
And $q(\boldsymbol{x}_{0},\boldsymbol{x}_{1})$ can be computed by the entropic optimal transport.
Unbalanced Wasserstein–Fisher–Rao metric ($g\neq 0,\sigma=0$) In this case, the model can account for the unbalanced dynamics, however, it can not account for stochastic dynamics. The cellular dynamics are also governed by the ODE model by $\mathrm{d}{\boldsymbol{x}}_{t}=\boldsymbol{b}({\boldsymbol{x}}_{t},t)\mathrm{d}t$. Then, one can use the dynamical unbalanced optimal transport to model these dynamics, also known as Wasserstein–Fisher–Rao metric Chizat et al. (2018a, b); Gangbo et al. (2019), which can be stated as
@@
Here, $\alpha$ denotes a hyperparameter that controls the weighting. It is also important to note that in this context, $\nu_{0}$ and $\nu_{1}$ do not necessarily correspond to normalized probability densities; rather, they generally represent mass densities.
@@
Recent works such as TrajectoryNet and TIGON utilize (16) to infer unbalanced dynamics from scRNA-seq data Sha et al. (2024); Tong et al. (2023). To derive a CNF solver for (16), TIGON Sha et al. (2024) observes that along the characteristic line $\frac{\mathrm{d}\boldsymbol{x}}{\mathrm{d}t}=\boldsymbol{b}(\boldsymbol{x},t)$, one has
$\int_{0}^{1}\int_{\mathbb{R}^{d}}f(\boldsymbol{x},t)p(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t=\mathbb{E}_{\boldsymbol{x_{0}}\sim p_{0}}\int_{0}^{1}f(\boldsymbol{x},t)e^{\int_{0}^{t}g(\boldsymbol{x},s)\mathrm{d}s}\mathrm{d}t$
@@
and $\frac{\mathrm{d}(\ln p)}{\mathrm{d}t}=g-\nabla\cdot\boldsymbol{v}$. This can make the computation of both energy loss and reconstruction loss in high dimensional space tractable. Therefore, one can parameterize $\boldsymbol{b}(\boldsymbol{x},t)$ and $g(\boldsymbol{x},t)$ using neural networks, respectively, and train them by minimizing the overall loss.
@@
Regularized Unbalanced Optimal Transport ($g\neq 0,\sigma\neq 0$) In this case, the model can account for both the unbalanced and stochastic dynamics. The cellular dynamics are governed by the SDE model $\mathrm{d}{\boldsymbol{x}}_{t}=\boldsymbol{b}\left({\boldsymbol{x}}_{t},t\right)\mathrm{d}t+\sigma(t)\boldsymbol{I}\mathrm{d}\boldsymbol{w}_{t}$. We can use the regularized unbalanced optimal transport to model it Zhang et al. (2025); Baradat and Lavenant (2021). It can be viewed as an unbalanced relaxation of the dynamic formulation of the Schrödinger bridge problem. Consider
$\inf_{\left(p,\boldsymbol{b},g\right)}\int_{0}^{1}\int_{\mathbb{R}^{d}}\frac{1}{2}\left\|\boldsymbol{b}(\boldsymbol{x},t)\right\|_{2}^{2}{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t+\int_{0}^{1}\int_{\mathbb{R}^{d}}\alpha\Psi\left(g(\boldsymbol{x},t)\right){p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t,$ (17)
@@
where $\Psi:\mathbb{R}\rightarrow[0,+\infty]$ corresponds to the growth penalty function, the infimum is taken over all pairs $\left(p,\boldsymbol{b}\right)$ such that $p(\cdot,0)=$ $\nu_{0},p(\cdot,1)=\nu_{1},p(\boldsymbol{x},t)$ absolutely continuous, and
@@
$\partial_{t}p=-\nabla_{\boldsymbol{x}}\cdot\left(p\boldsymbol{b}\right)+\frac{1}{2}\nabla_{\boldsymbol{x}}^{2}:\left(\sigma^{2}(t)\boldsymbol{I}p\right)+gp$ (18)
with vanishing boundary condition: $\lim_{|\boldsymbol{x}|\rightarrow\infty}p(\boldsymbol{x},t)=0$.
[…]
@@
$\inf_{\left(p,\boldsymbol{v},g\right)}\int_{0}^{1}\int_{\mathbb{R}^{d}}\left[\frac{1}{2}\left\|\boldsymbol{v}(\boldsymbol{x},t)\right\|_{2}^{2}+\frac{\sigma^{4}(t)}{8}\left\|\nabla_{\boldsymbol{x}}\log p\right\|_{2}^{2}-\frac{\sigma^{2}(t)}{2}\left(1+\log p\right)g+\alpha\Psi\left(g\right)\right]{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t,$ (19)